# PULASki: Learning inter-rater variability using statistical distances to improve probabilistic segmentation

### v1

@@ 2.5 Stochastic Segmentation Networks (SSN)
Stochastic Segmentation Networks Monteiro2020 were proposed to model aleatoric uncertainty primarily due to integrader variability. Unlike the Probabilistic U-Net, the SSN can be more easily used with any neural network based segmentation model. The logits (i.e. inputs to the softmax) $\eta\in\mathbb{R}^{d^{2}.K\times 1}$, represented here as a flattened vector, are modelled as a Gaussian random variable whose mean and covariance are modelled by neural networks,
@@
where $\eta^{n,m}\sim N(\mu(x^{n}),\Sigma(x^{n}))$ as defined in 19. Note that 22 takes a similar form to 12 but without a regularisation term due to the lack of the Bayesian update. This may potentially lead to a less tractable optimisation problem that may have difficulty converging poggio2020complexity .
@@ 2.6 Monte Carlo Dropout
Monte Carlo Dropout (MCDO) is a widespread tool for approximating model uncertainty in neural networks by randomly removing layers during both training and prediction phases kendall2015bayesian ; gal2016dropout . In the context of the U-Net, MCDO can be implemented by adding spatial dropout layers tompson2015efficient after every encoding (pair of convolution layers with batch normalisation and ReLU activations, followed by average pooling) and decoding block (transposed convolution, followed by a pair of convolution layers wit batch normalisation and ReLU). The key idea behind MCDO is to randomly switch on and off neurons (i.e. set their weights to zero) according to a pre-specified probability. Running this procedure $M$ times allows one to generate $M$ segmentations. However, it is noteworthy that even though utilising the Dropout function is computationally inexpensive, there are a number of intrinsic shortcomings of MCDO, such as unreliable uncertainty quantification, subpar calibration, and inadequate out-of-distribution detection Fuchs2022 .
@@ 3 PULASki - Probabilistic Unet Loss Assessed through Statistical distances
Our main focus is to improve the capability of U-Net based image segmentation methods to capture predicted aleatoric uncertainty, primarily due to inter-rater variability. Inter-rater variability is a persistent issue in biomedical segmentation which arises primarily due to the low signal to noise ratio in the raw data and complexity of the features being segmented (e.g., small cerebral vessels). In the original formulation of the Probabilistic U-Net, the reconstruction term in the ELBO 10 can be seen as an empirical approximation of the expected cross-entropy,
@@
and $\mathcal{Y}$ is the set of all possible segmentations for a given image. The standard cross-entropy has several known shortcomings for class-imbalanced datasets and for multi-label classification problems abraham2019novel ; rezaei2020addressing ; tian2022striking . An empirical approximation of the cross-entropy between $p(y|x)$ and $p_{\theta}(y|z,x)$ when used to approximate the ELBO is indistinguishable from the standard ELBO (12) when multiple labels are treated as conditionally independent data. This makes it poorly suited for learning the conditional distribution $p(y|x)$.
@@
where $D(p,q)$ is an appropriate statistical distance between two probability distributions $p,q$. In practice, distances are usually approximated using the empirical distributions; Samples from $p_{\theta}(y|x,z)$ can be easily generated during training using the same operations as for the Probabilistic U-Net (see 13 - 15. An empirical approximation of $p(y|x)$ is available whenever there are multiple plausible segmentations per image. A schematic of the method can be found in Figure 2. Our proposed method bears some resemblance to wasserstein autoencoders Tolstikhin2018 , although our primary focus is on varying the reconstruction loss for conditioned generation whereas their focus is on alternative regularisation terms.
Crucially, all statistical distances considered here are differentiable, so that they can be used as loss-functions in standard gradient-based optimisation methods such as ADAM. We focus on distances that are well known in the image analysis literature, as described below.
@@ 3.1 Frechet Inception Distance (FID)
The FID Heusel2017 is a pseudometric based on the Wasserstein-2 distance and is widely used particularly in training GANs for image generation. Computing the Wasserstein-2 distance between two arbitrary empirical distributions can be computationally prohibitive in DL applications which require repeated evaluations during training. Calculation of the FID requires approximating the distributions by a Gaussian, which then allows one to exploit the existence of a closed form solution to the Wasserstein-2 distance for Gaussian distributions, making it fast to implement. It provides a reasonable approximation for distributions whose first two moments exist, although it can produce biased estimates Binkowski2018 .
@@
where $\hat{\mu}\in\mathbb{R}^{v}$ and $\hat{\Sigma}\in\mathbb{R}^{v\times v}$ are the sample mean and covariance respectively of $\{a^{m}\}_{m=1:M}$ where $a^{m}$ denotes a flattened (i.e. vectorised) version of the matrix $\eta^{m}$ as defined in 15 of length $v=d\times d\times K$. Likewise, $\mu$ and $\Sigma$ denote the sample mean and covariance respectively of the final activations (i.e. logits) from an Inception Net V3 szegedy2016rethinking which takes the plausible segmentations rather than raw images as inputs. Since the inception net is only trained on 2D images, we only make use of the FID for the experiments involving 2D images.
@@ 3.2 Hausdorff Divergence
The Hausdorff distance is a way of measuring distances between subsets of metric spaces (rather than measure spaces). It has traditionally been used in computer vision to measure the degree of similarity between two objects, such as image segmentations, by measuring distances between points of the two sets. The traditional Hausdorff distance can be lifted to measuring statistical distance between probability measures. In the case of discrete measures $\alpha,\beta$ on $\mathbb{R}^{d}$ given as weighted point clouds i.e.
$\alpha=\sum_{i=1}^{N}\alpha_{i}\delta_{x_{i}};\quad\beta=\sum_{i=1}^{N}\beta_{i}\delta_{y_{i}};$ (28)
where $\delta_{x}$ is the Dirac delta function centred at $x$, the Hausdorff divergence is defined (as per definition A.2 in Feydy2020 ) as
$d_{H}(\alpha,\beta):=\frac{1}{2}\langle\alpha-\beta,\nabla F_{\epsilon}(\alpha)-\nabla F_{\epsilon}(\beta)\rangle.$ (29)
@@
$F_{\epsilon}(\alpha):=-\frac{1}{2}\mathrm{OT}_{\epsilon}(\alpha,\alpha)$ (30)
@@
$OT_{\epsilon}(\alpha,\beta)=\min_{\pi\in M^{+}(X\times X)}\left\langle\pi,C(x,y)\right\rangle+\epsilon KL(\pi||\alpha\otimes\beta)$ (31)
@@
The above is more specifically referred to as the $\epsilon$-Hausdorff divergence, as it is based on $OT_{\epsilon}$ which only produces an $\epsilon$-approximate solution to the exact OT problem. The $\epsilon$-Hausdorff divergence is generally cheaper to compute than the closely related $\epsilon$-Sinkhorn divergence or debiased Sinkhorn divergence (see Section 3.3) as it relies on solving the symmetric OT problem $OT_{\epsilon}(\alpha,\alpha)$ rather than $OT_{\epsilon}(\alpha,\beta)$.
@@
A more accurate but computationally expensive approximation of the $p$-Wasserstein distance is the so-called de-biased Sinkhorn divergence Ramdas2017
$S_{\epsilon}(\alpha,\beta):=\text{OT}_{\epsilon}(\alpha,\beta)-\frac{1}{2}\text{OT}_{\epsilon}(\alpha,\alpha)-\frac{1}{2}\text{OT}_{\epsilon}(\beta,\beta)$ (32)
where $\alpha,\beta$ are any positive measures and $\text{OT}_{\epsilon}$ is the entropic-regularised form of the optimal transport loss 31. This divergence has been used in training generative models (See e.g. genevay18a ). We utilise the method proposed in Feydy2020 which makes use of a multiscale Sinkhorn algorithm to compute $S_{\epsilon}(\alpha,\beta)$.
@@ 4.1 Datasets and Labels
To be able to validate the proposed approach, the OpenNeuro’s ‘StudyForrest’ dataset which contains 20 subjects was used for vessel segmentation. It contains 7T MRA-ToF full brain volumetric images with [480 × 640 × 163] matrix size in NIFTI format recorded with a Siemens MR scanner at 7-Tesla using a 3D Multiple Overlapping Thin Slab Acquisition Time of Flight angiography sequence (0.3 mm isotropic voxel size) hausler2021studyforrest . Training the network requires multiple plausible segmentations. For the ‘StudyForrest’ dataset, multiple manual labels did not exist. Therefore, ten synthetic plausible segmentations per case have been generated using the Frangi filter frangi1998 . The Frangi scale range has been set to $\sigma$ $\sim$ [1 8] and a step size between the sigmas of 2 was chosen. Following parameters, known as vesselness constants, have been sampled from truncated normal distributions: (i) $\mathrm{Frangi}_{\alpha}$ $\sim$ TN(0.5 1) determines if its a line (vessel) or plane like structure; (ii) $\mathrm{Frangi}_{\beta}$ $\sim$ TN(0.5 1) determines the deviation from a blob like structure; (iii) $\mathrm{Frangi}_{C}$ $\sim$ TN(500 100) gives a threshold between eigenvalues of noise and vessel structures. The resulting ”inter-rater” agreement of the generated dataset was calculated using $K\alpha$ (discussed later in Sec. 4.3) and has a value of 0.40.
For further validating the proposed method, 3T FLAIR MRIs from the publicly available multiple sclerosis (MS) segmentation dataset, that was used for MICCAI 2016 challenge commowick2021multiple , containing MRI scans of 53 patients and their corresponding manual segmentations of the lesions by seven experts, was utilised. After thorough examination of expert voting, one subject was excluded due to significant discrepancies in expert annotations. In this case, two graders determined that no voxels represented an MS lesion, four experts identified only a few voxels as MS tissue, and one grader marked nearly the entire brain. Accurate identification of multiple sclerosis lesions is difficult due to variability in lesion location, size and shape – which is also reflected in the resulting very low inter-rater agreement. The inter-rater agreement was assessed using $K\alpha$ for the MS dataset, resulting in an agreement of 0.60. However, when considering only brain regions where at least one grader noted an MS occurrence, the alpha decreased to 0.28.
@@ 4.2 Implementation
The PULASki method was implemented by extending the pipeline  [^1] DS6 Pipeline: https://github.com/soumickmj/DS6 available from DS6 Chatterjee2020 . It is based on PyTorch paszke2019pytorch and used TorchIO perez2021torchio for creating 3D patches or extracting 2D slices from the available 3D volumes. All the statistical distances used in PULASKi (discussed in Sec. 3) were calculated using the GeomLoss [^2] GeomLoss on GitHub: https://github.com/jeanfeydy/geomloss feydy2019interpolating . The weights and biases of all the models (proposed and baselines, including their sub-models – e.g., prior net), were initialised by sampling a uniform distribution $\mathcal{U}(-\sqrt{k},\sqrt{k})$, where $k=\frac{\text{ groups }}{C_{\text{in }}*\prod_{i=0}^{1}\text{ kernel\_size }[i]}$ for the convolutional layers and $k=\frac{1}{in\_features}$ for the linear layers.
@@
The baseline probabilistic U-Net experiments were performed by adapting the PyTorch implementation [^3] PyTorch implementation of Probabilistic UNet baseline: https://github.com/jenspetersen/probabilistic-unet available from petersen2019deep . As the loss function for the baseline probabilistic UNet, as well as for the MCDO UNet, Focal Tversky Loss (FTL) abraham2019novel was employed. Cross-entropy (CE) loss used in the original probabilisitic UNet work was replaced with FTL as CE might not be optimal when there is a big imbalance between the positive (1) and negative (0) classes in the image mask as CE treats them both equally – a common scenario in medical image segmentation tasks rajaraman2021novel ; tian2022striking . A balanced version of CE loss (weighted CE) might combat this issue tian2022striking . Dice loss, on the other hand, is a commonly used loss function in medical image segmentation tasks, providing a better trade-off between precision and recall abraham2019novel . FTL, designed specifically for medical image segmentation tasks, achieves an even better trade-off between precision and recall when training on small structures abraham2019novel , and has also been shown to outperform CE, weighted CE, and Dice rajaraman2021novel .
@@
The SSN method was performed using the DeepMedic convolutional neural network KAMNITSAS201761 as backbone, as was done in the original paper Monteiro2020 . DeepMedic is a dual pathway, 11 layers deep, three-dimensional convolutional neural network, that was developed for brain lesion segmentation. The network has two parallel convolutional pathways that process the input images at multiple scales simultaneously, capturing both local and larger contextual information, and employs a novel training scheme that alleviates the class-imbalance problem and enables dense training on 3D image segments. Contrary to U-Nets, which condense and then expand the input to produce the final output, DeepMedic employs a fully convolutional network that processes the input at different resolutions through two parallel pathways, then uses a series of convolutional and pooling layers to extract features at different scales, and finally combines these features through concatenation and further convolutional layers to provide the output.
The final baseline model, MH-VI U-Net Fuchs2022 , was implemented by adapting the official code [^4] MH-VI U-Net Official code: https://github.com/MECLabTUDA/VIMH. All parameters, including four heads with a $\lambda$ of $0.5$, were kept the same as the original implementation.
@@
The models were trained by optimising the respective loss function for that particular model (as discussed earlier) using the Adam optimiser for 500 epochs. It is worth mentioning that all the models reached convergence before the pre-set number of epochs and the model state with the lowest validation loss was chosen for final evaluation. To reduce the training time per epoch, a limit of randomly chosen patch/slice was chosen with a uniform probability from all available samples, for each experiment – 1500 and 4000, for 2D and 3D, respectively for the task of vessel segmentation, and 10000 for the task of MS segmentation. Increasing the limit will not have any impact on the results as all models converged before the 500th epoch – rather, it will increase the time to finish 500 epochs, while converging after an earlier epoch. Trainings were performed using Nvidia 2080 TI, V100, A6000 GPUs – chosen depending upon the computational demands of each model. The complete implementation of this project (PULASki method, as well as the baselines used here) is publicly available on GitHub [^5] PULASki on GitHub: https://github.com/soumickmj/PULASki.

### v2

@@ 2.7 Stochastic Segmentation Networks (SSN)
Stochastic Segmentation Networks [14] were proposed to model aleatoric uncertainty primarily due to integrader variability. Unlike the Probabilistic U-Net, the SSN can be more easily used with any neural network based segmentation model. The logits (i.e. inputs to the softmax) $\eta\in\mathbb{R}^{d^{2}.K\times 1}$, represented here as a flattened vector, are modelled as a Gaussian random variable whose mean and covariance are modelled by neural networks,
@@
where $\eta^{n,m}\sim N(\mu(x^{n}),\Sigma(x^{n}))$ as defined in 32. Note that 35 takes a similar form to 12 but without a regularisation term due to the lack of the Bayesian update. This may potentially lead to a less tractable optimisation problem that may have difficulty converging [37].
@@ 2.8 Monte Carlo Dropout
Monte Carlo Dropout (MCDO) is a widespread tool for approximating model uncertainty in neural networks by randomly removing layers during both training and prediction phases [38, 39]. In the context of the U-Net, MCDO can be implemented by adding spatial dropout layers [23] after every encoding (pair of convolution layers with batch normalisation and ReLU activations, followed by average pooling) and decoding block (transposed convolution, followed by a pair of convolution layers wit batch normalisation and ReLU). The key idea behind MCDO is to randomly switch on and off neurons (i.e. set their weights to zero) according to a pre-specified probability. Running this procedure $M$ times allows one to generate $M$ segmentations. However, it is noteworthy that even though utilising the Dropout function is computationally inexpensive, there are a number of intrinsic shortcomings of MCDO, such as unreliable uncertainty quantification, subpar calibration, and inadequate out-of-distribution detection [26].
@@ 3 PULASki - Probabilistic Unet Loss Assessed through Statistical distances
Our main focus is to improve the capability of U-Net based image segmentation methods to capture predicted aleatoric uncertainty, primarily due to inter-rater variability. Inter-rater variability is a persistent challenge in biomedical segmentation, often driven by factors such as limited image clarity (e.g., lower signal-to-noise ratio, presence of artefacts) and the inherent complexity of the structures being segmented (e.g., small cerebral vessels). In the original formulation of the Probabilistic U-Net, the reconstruction term in the ELBO 10 can be seen as an empirical approximation of the expected cross-entropy,
@@
and $\mathcal{Y}$ is the set of all possible segmentations for a given image. The standard cross-entropy has several known shortcomings for class-imbalanced datasets and for multi-label classification problems [13, 27, 28]. An empirical approximation of the cross-entropy between $p(y|x)$ and $p_{\theta}(y|z,x)$ when used to approximate the ELBO is indistinguishable from the standard ELBO (12) when multiple labels are treated as conditionally independent data. This makes it poorly suited for learning the conditional distribution $p(y|x)$.
@@
where $D(p,q)$ is an appropriate statistical distance between two probability distributions $p,q$. In practice, distances are usually approximated using the empirical distributions; Samples from $p_{\theta}(y|x,z)$ can be easily generated during training by applying the same latent vector operations as in the Prob U-net $M$ times (see 13 - 15). An empirical approximation of $p(y|x)$ is available whenever there are multiple plausible segmentations per image. A schematic of the method can be found in Figure 2. Our proposed method bears some resemblance to wasserstein autoencoders [40], although our primary focus is on varying the reconstruction loss for conditioned generation whereas their focus is on alternative regularisation terms (i.e. the KL term).
Algorithm 1 PULASki Training
Training dataset $\{(x_{i},\{y_{i}^{j}\}_{j=1}^{M})\}_{i=1}^{N}$ with $N$ images and $M$ ground-truth segmentations per image
Statistical distance function $D(P,Q)$
Weight for KL divergence term, $\beta$
Trained model parameters
for each training iteration do
     Sample a minibatch of input images $\{x_{b}\}_{b=1}^{B}$ and corresponding ground-truth segmentations $\{\{y_{b}^{j}\}_{j=1}^{M}\}_{b=1}^{B}$
     for each training iteration do
         Sample a minibatch of input images $\{x_{b}\}_{b=1}^{B}$ and corresponding ground-truth segmentations $\{\{y_{b}^{j}\}_{j=1}^{M}\}_{b=1}^{B}$
         for each image $x_{b}$ in the minibatch do
              Encode $x_{b}$ to obtain prior parameters: $p_{\theta}(z|x_{b})$
              Encode $x_{b}$ and each $y_{b}^{j}$ to obtain posterior parameters: $q_{\phi}(z|x_{b},y_{b}^{j})$
              Sample $M$ latent variables $\{z_{b}^{j}\}_{j=1}^{M}$ from the posterior distributions $q_{\phi}(z|x_{b},y_{b}^{j})$
              Generate $M$ predicted segmentations $\{\hat{y}_{b}^{j}\}_{j=1}^{M}$ using the decoder: $\hat{y}_{b}^{j}=\text{Decoder}(x_{b},z_{b}^{j})$
              Calculate statistical distance loss: $L_{\text{StatDist}}=D(\{\hat{y}_{b}^{j}\}_{j=1}^{M},\{y_{b}^{j}\}_{j=1}^{M})$
              Calculate KL divergence for each pair: $L_{\text{KL}}=\frac{1}{M}\sum_{j=1}^{M}\text{KL}(q_{\phi}(z|x_{b},y_{b}^{j})\parallel p_{\theta}(z|x_{b}))$
              Compute total loss for $x_{b}$: $L_{b}=L_{\text{StatDist}}+\beta\cdot L_{\text{KL}}$
         end for
         Compute average loss over the minibatch: $L=\frac{1}{B}\sum_{b=1}^{B}L_{b}$
         Update model parameters to minimise $L$ using backpropagation
     end for
     Compute average loss over the minibatch: $L=\frac{1}{B}\sum_{b=1}^{B}L_{b}$
     Update model parameters to minimise $L$ using backpropagation
end for
Crucially, all statistical distances considered here are differentiable, so that they can be used as loss-functions in standard gradient-based optimisation methods such as ADAM. Furthermore, distances focusing on both distributional quality as well as local geometric features are chosen. Broadly speaking, FID and Sinkhorn divergence are used to measure global distributional differences between samples of images, whereas the Hausdorff divergence is primarily used to measure differences between sets of points, thereby emphasising differences in local geometric structure. Therefore, both FID and Sinkhorn are more widely used in the image generation literature (e.g. GANs and diffusion modelling) whereas Hausdorff divergence is used in used in shape analysis and object detection in the medical imaging context. The Hausdorff divergence is much more sensitive to differences in local image features such as edges and contours, making it a useful tool for comparing segmentations. The considered distances (FID, Hausdorff and Sinkhorn) are described in further detail below.
@@ 3.1 Frechet Inception Distance (FID)
The FID [41] is a pseudometric based on the Wasserstein-2 distance and is widely used particularly in training GANs for image generation. Computing the Wasserstein-2 distance between two arbitrary empirical distributions can be computationally prohibitive in DL applications which require repeated evaluations during training. Calculation of the FID requires approximating the distributions by a Gaussian, which then allows one to exploit the existence of a closed form solution to the Wasserstein-2 distance for Gaussian distributions, making it fast to implement. It provides a reasonable approximation for distributions whose first two moments exist, although it can produce biased estimates [42].
@@
where $\hat{\mu}\in\mathbb{R}^{v}$ and $\hat{\Sigma}\in\mathbb{R}^{v\times v}$ are the sample mean and covariance respectively of $\{a^{m}\}_{m=1:M}$ where $a^{m}$ denotes a flattened (i.e. vectorised) version of the matrix $\eta^{m}$ as defined in 15 of length $v=d\times d\times K$. Likewise, $\mu$ and $\Sigma$ denote the sample mean and covariance respectively of the final activations (i.e. logits) from an Inception Net V3 [43] which takes the plausible segmentations rather than raw images as inputs. FID, employing Inception Net V3 trained on ImageNet, provides a viable metric for medical images by quantifying the statistical alignment of low-level visual features (e.g., edges, textures) that are inherently universal across image domains. Since FID evaluates the distributional similarity of these domain-agnostic patterns, rather than relying on semantic or medical specificity, it offers a pragmatic solution for assessing generative models in medical imaging contexts. Since the inception net is only trained on 2D images, we only make use of the FID for the experiments involving 2D images.
@@ 3.2 Hausdorff Divergence
The Hausdorff distance is a way of measuring distances between subsets of metric spaces (rather than measure spaces). It has traditionally been used in computer vision to measure the degree of similarity between two objects, such as image segmentations, by measuring distances between points of the two sets. The traditional Hausdorff distance can be lifted to measuring statistical distance between probability measures. In the case of discrete measures $\vartheta,\xi$ on $\mathbb{R}^{d}$ given as weighted point clouds i.e.
$\vartheta=\sum_{i=1}^{N}\vartheta_{i}\delta_{x_{i}};\quad\xi=\sum_{i=1}^{N}\xi_{i}\delta_{y_{i}};$ (41)
where $\delta_{x}$ is the Dirac delta function centred at $x$, the Hausdorff divergence is defined (as per definition A.2 in [44]) as
$d_{H}(\vartheta,\xi):=\frac{1}{2}\langle\vartheta-\xi,\nabla F_{\epsilon}(\vartheta)-\nabla F_{\epsilon}(\xi)\rangle.$ (42)
@@
$F_{\epsilon}(\vartheta):=-\frac{1}{2}\mathrm{OT}_{\epsilon}(\vartheta,\vartheta)$ (43)
@@
$OT_{\epsilon}(\vartheta,\xi)=\min_{\pi\in M^{+}(X\times X)}\left\langle\pi,C(x,y)\right\rangle+\epsilon KL(\pi||\vartheta\otimes\xi)$ (44)
@@
The above is more specifically referred to as the $\epsilon$-Hausdorff divergence, as it is based on $OT_{\epsilon}$ which only produces an $\epsilon$-approximate solution to the exact OT problem. The $\epsilon$-Hausdorff divergence is generally cheaper to compute than the closely related $\epsilon$-Sinkhorn divergence or debiased Sinkhorn divergence (see Section 3.3) as it relies on solving the symmetric OT problem $OT_{\epsilon}(\vartheta,\vartheta)$ rather than $OT_{\epsilon}(\vartheta,\xi)$.
@@
A more accurate but computationally expensive approximation of the $p$-Wasserstein distance is the so-called de-biased Sinkhorn divergence [45]
$S_{\epsilon}(\vartheta,\xi):=\text{OT}_{\epsilon}(\vartheta,\xi)-\frac{1}{2}\text{OT}_{\epsilon}(\vartheta,\vartheta)-\frac{1}{2}\text{OT}_{\epsilon}(\xi,\xi)$ (45)
where $\vartheta,\xi$ are any positive measures and $\text{OT}_{\epsilon}$ is the entropic-regularised form of the optimal transport loss 44. This divergence has been used in training generative models (See e.g. [46] ). We utilise the method proposed in [44] which makes use of a multiscale Sinkhorn algorithm to compute $S_{\epsilon}(\vartheta,\xi)$.
@@ 4.1 Datasets and Labels
To be able to validate the proposed approach, the OpenNeuro’s ‘StudyForrest’ dataset which contains 20 subjects was used for vessel segmentation. It contains 7T MRA-ToF full brain volumetric images with [480 × 640 × 163] matrix size in NIFTI format recorded with a Siemens MR scanner at 7-Tesla using a 3D Multiple Overlapping Thin Slab Acquisition Time of Flight angiography sequence (0.3 mm isotropic voxel size) [47]. Training the network requires multiple plausible segmentations. For the ‘StudyForrest’ dataset, multiple manual labels did not exist. Therefore, ten synthetic plausible segmentations per case have been generated using the Frangi filter [48]. The Frangi scale range has been set to $\sigma$ $\sim$ [1 8] and a step size between the sigmas of 2 was chosen. Following parameters, known as vesselness constants, have been sampled from truncated normal distributions: (i) $\mathrm{Frangi}_{\alpha}$ $\sim$ TN(0.5 1) determines if its a line (vessel) or plane like structure; (ii) $\mathrm{Frangi}_{\beta}$ $\sim$ TN(0.5 1) determines the deviation from a blob like structure; (iii) $\mathrm{Frangi}_{C}$ $\sim$ TN(500 100) gives a threshold between eigenvalues of noise and vessel structures. The resulting ”inter-rater” agreement of the generated dataset was calculated using $K\alpha$ (discussed later in Sec. 4.3) and has a value of 0.40. The labels generated during this research are made available publicly as part of the SMILE-UHURA dataset [49].
For further validating the proposed method, 3T FLAIR MRIs from the publicly available multiple sclerosis (MS) segmentation dataset, that was used for MICCAI 2016 challenge [50], containing MRI scans of 53 patients and their corresponding manual segmentations of the lesions by seven experts, was utilised. After thorough examination of expert voting, one subject was excluded due to significant discrepancies in expert annotations. In this case, two graders determined that no voxels represented an MS lesion, four experts identified only a few voxels as MS tissue, and one grader marked nearly the entire brain. Accurate identification of multiple sclerosis lesions is difficult due to variability in lesion location, size and shape – which is also reflected in the resulting very low inter-rater agreement. The inter-rater agreement was assessed using $K\alpha$ for the MS dataset, resulting in an agreement of 0.60. However, when considering only brain regions where at least one grader noted an MS occurrence, the alpha decreased to 0.28.
@@ 4.2 Implementation
The PULASki method was implemented by extending the pipeline  [^1] DS6 Pipeline: https://github.com/soumickmj/DS6 available from DS6 [12]. It is based on PyTorch [51] and used TorchIO [52] for creating 3D patches or extracting 2D slices from the available 3D volumes. All the statistical distances used in PULASKi (discussed in Sec. 3) were calculated using the GeomLoss [^2] GeomLoss on GitHub: https://github.com/jeanfeydy/geomloss [53]. The weights and biases of all the models (proposed and baselines, including their sub-models – e.g., prior net), were initialised by sampling a uniform distribution $\mathcal{U}(-\sqrt{k},\sqrt{k})$, where $k=\frac{\text{ groups }}{C_{\text{in }}*\prod_{i=0}^{1}\text{ kernel\_size }[i]}$ for the convolutional layers and $k=\frac{1}{in\_features}$ for the linear layers.
@@
The baseline probabilistic U-Net experiments were performed by adapting the PyTorch implementation [^3] PyTorch implementation of Probabilistic UNet baseline: https://github.com/jenspetersen/probabilistic-unet available from [54]. As the loss function for the baseline probabilistic UNet, as well as for the MCDO UNet, Focal Tversky Loss (FTL) [13] was employed. Cross-entropy (CE) loss used in the original probabilisitic UNet work was replaced with FTL as CE might not be optimal when there is a big imbalance between the positive (1) and negative (0) classes in the image mask as CE treats them both equally – a common scenario in medical image segmentation tasks [55, 28]. A balanced version of CE loss (weighted CE) might combat this issue [28]. Dice loss, on the other hand, is a commonly used loss function in medical image segmentation tasks, providing a better trade-off between precision and recall [13]. FTL, designed specifically for medical image segmentation tasks, achieves an even better trade-off between precision and recall when training on small structures [13], and has also been shown to outperform CE, weighted CE, and Dice [55].
@@
The SSN method was performed using the DeepMedic convolutional neural network [56] as backbone, as was done in the original paper [14]. DeepMedic is a dual pathway, 11 layers deep, three-dimensional convolutional neural network, that was developed for brain lesion segmentation. The network has two parallel convolutional pathways that process the input images at multiple scales simultaneously, capturing both local and larger contextual information, and employs a novel training scheme that alleviates the class-imbalance problem and enables dense training on 3D image segments. Contrary to U-Nets, which condense and then expand the input to produce the final output, DeepMedic employs a fully convolutional network that processes the input at different resolutions through two parallel pathways, then uses a series of convolutional and pooling layers to extract features at different scales, and finally combines these features through concatenation and further convolutional layers to provide the output.
MH-VI U-Net [26], was implemented by adapting the official code [^4] MH-VI U-Net official code: https://github.com/MECLabTUDA/VIMH. All parameters, including four heads with a $\lambda$ of $0.5$, were kept the same as the original implementation. Similarly, D-Persona [35] and CIMD [36] baseline models were also implemented adapting their respective official codes  [^5] D-Persona official code: https://github.com/ycwu1997/D-Persona, [^6] CIMD official code: https://github.com/aimansnigdha/Ambiguous-Medical-Image-Segmentation-using-Diffusion-Models, and all default parameters were utilised. During inference with CIMD, the number of timesteps $T$ was reduced from 1000 to 100 to minimise the time required for inference, as it is significantly longer than that of other methods. All the probabilistic U-Nets (including the PULASki methods) required approximately 10 and 30 seconds per volume for MS and vessel segmentation, respectively. By contrast, the MH-VI U-Net (the second slowest baseline) required approximately 30 seconds and 6.5 minutes, respectively, for the same tasks. On the same hardware, CIMD, even with the reduced $T=100$, required over 1 hour for MS segmentation and more than 6 hours for vessel segmentation (with $T=1000$, it took more than 10 and 60 hours, respectively). Furthermore, as no stopping criterion for CIMD was specified in either the paper or the code, early stopping with a patience of 5 epochs for no change in the loss was implemented during model training.
@@
The models were trained by optimising the respective loss function for that particular model (as discussed earlier) using the Adam optimiser for 500 epochs. It is worth mentioning that all the models reached convergence before the pre-set number of epochs and the model state with the lowest validation loss was chosen for final evaluation. To reduce the training time per epoch, a limit of randomly chosen patch/slice was chosen with a uniform probability from all available samples, for each experiment – 1500 and 4000, for 2D and 3D, respectively for the task of vessel segmentation, and 10000 for the task of MS segmentation. Increasing the limit will not have any impact on the results as all models converged before the 500th epoch – rather, it will increase the time to finish 500 epochs, while converging after an earlier epoch. Trainings were performed using Nvidia 2080 TI, V100, A6000 GPUs – chosen depending upon the computational demands of each model. Automatic mixed precision was employed for all the models, to reduce GPU memory requirements. The complete implementation of this project (PULASki method, as well as the baselines used here) is publicly available on GitHub [^7] PULASki on GitHub: https://github.com/soumickmj/PULASki, while the trained model weights are available on Hugging Face  [^8] PULASki weights on Hugging Face: https://huggingface.co/collections/soumickmj/pulaski-66d9d35dfef91c84d140de8d.